### v1

# A Survey of Cognitive Distortion Detection and Classification in NLP

Abstract
As interest grows in the application of natural language processing techniques to mental health, a growing body of work explores the automatic detection and classification of cognitive distortions. cognitive distortions are habitual patterns of negatively biased or flawed thinking that distort how people perceive events, judge themselves, and react to the world around them. Identifying and addressing them is an important part of therapy. Despite its momentum, the field remains fragmented, with inconsistencies in cognitive distortion taxonomies, task formulations, and evaluation practices. This survey reviews 38 studies spanning two decades, providing a structured overview of datasets, modelling approaches, and evaluation strategies. We provide a consolidated cognitive distortion taxonomy reference, summarise common task setups, and highlight open challenges to support more coherent and reproducible research in this emerging area.
A Survey of Cognitive Distortion Detection and Classification in NLP
Archie Sage, Jeroen Keppens, Helen Yannakoudakis Department of Informatics, King’s College London {archie.sage, jeroen.keppens, helen.yannakoudakis}@kcl.ac.uk
1 Introduction
Cognitive distortions are habitual patterns of negatively biased or logically flawed thinking that distort how people perceive events, judge themselves, and react to the world around them. These distortions play a central role in emotional distress and are a core target of evidence-based psychological interventions such as cognitive behavioural therapy (Beck, 1963; Burns, 1999).
Common examples include Catastrophising (‘I let them down, so they’ll never trust me again’), Mind Reading (‘They haven’t replied, so they must be angry at me’), and All or Nothing Thinking (‘If I don’t get this right the first time, I’m a complete failure’). [^1] A complete list of cognitive distortion definitions, examples, and synonyms is provided in Table 3, with additional psychological background in Appendix B. These patterns often appear intuitive or harmless at first glance, but they have been shown to maintain and exacerbate conditions like depression, anxiety, and post-traumatic stress disorder. In therapeutic settings, recognising and reframing such distortions is a core goal of cognitive behavioural therapy. Crucially, these distortions are primarily expressed through language, making them well-suited to computational modelling. Recent work in natural language processing has begun to explore the automatic detection and classification of cognitive distortions, with applications ranging from clinical decision support tools to mental health chatbots, journaling tools, and triage systems. Studies have shown that incorporating cognitive distortion-level features can improve outcomes in related tasks such as depression detection (Wang et al., 2023b), complementing more traditional sentiment or topic based approaches. By identifying distorted cognitive patterns in everyday text, natural language processing systems may support more timely, personalised, and psychologically-informed interventions.
Despite rapid growth in the field, the literature remains fragmented. Computational approaches use inconsistent taxonomies for defining cognitive distortions, making it difficult to compare findings across studies. Task formulations (e.g., detection vs classification, single-label vs multi-label) vary widely, often reflecting implicit assumptions that shape evaluation and outcomes. Benchmarks are scarce, metrics inconsistently applied, and variations in domain and dataset usage further complicate comparisons, making it hard to pinpoint gaps or establish best practices.
[…]
Table 1: Consolidated cognitive distortion categories and their inclusion across natural language processing papers in our survey that address the cognitive distortion classification task, either directly or conceptually. Where applicable, papers are grouped by the cognitive distortion dataset they rely on, reflecting the underlying taxonomy of those datasets. Definitions of cognitive distortion categories can be found in Table 3.
[…]
To our knowledge, the most comparable prior survey is by Suputra et al. (2023), which examined 12 studies and provided an initial synthesis of modelling approaches used at the time. Our work expands on this by covering 38 publications, including recent preprints, and offering a structured overview of the computational landscape for cognitive distortion detection and classification. This survey aims to provide a clear and practical entry point for researchers engaged in the growing field of cognitive distortion detection and classification from a computational perspective. Our focus is firmly on how computational methods approach these tasks, without seeking to redefine or adjudicate psychological constructs themselves. The contributions in this paper are threefold. (i) It provides a consolidated reference of the cognitive distortion taxonomies used across computational studies, highlighting inconsistencies and listing common synonyms (Tables 1 and 3). (ii) It defines and analyses task setups, datasets, computational methods, and approaches to performance evaluation, highlighting key patterns and gaps (§2, §4, §5, §6). (iii) It identifies open challenges and proposes concrete directions to guide future research (§7). In doing so, this paper aims to enable more consistent, comparable, and reproducible work in this emerging field.
2 Task Definitions
At its core, computational work on cognitive distortions involves a basic classification task: determining whether a given text reflects distorted thinking, and if so, identifying the specific type(s) of distortion. However, a number of different versions of this definition have been adopted in distinct groups of studies, inhibiting direct comparison between them. Some publications combine the classification task with additional tasks that may or may not inform classification. This section reviews how existing research defines cognitive distortion tasks, highlights key differences, and links these to their clinical foundations.
2.1 Binary Classification (Detection)
The simplest way to frame the computational task is to ask whether a given text span contains any instance of distorted thinking. This is often called detection and is usually treated as a binary classification problem (Distorted vs Undistorted). Methodologically, detection is a basic case of classification with just two labels - the approach remains the same, only the label set is coarser. This binary framing reflects early clinical aims, where simply noticing the presence of distorted thinking is an important first step before exploring the person’s thoughts in more detail during therapy. Even so, most computational studies go further, aiming to identify specific cognitive distortion types.
2.2 Single-label vs Multi-label Classification
While detection addresses the presence of distorted thinking, classification aims to specify which type(s) of cognitive distortion are present in a given text. Here, task definitions diverge based on assumptions around label cardinality:
Single-label classification.
This assumes that each text span reflects exactly one type of distortion. This simplifying assumption is often motivated by practical constraints, such as dataset design or the brevity of inputs (e.g., tweets, short user queries). However, it neglects the fact that distorted thoughts frequently exhibit multiple overlapping cognitive distortion categories, particularly in longer or more detailed texts.
Multi-label classification.
This allows a text to be assigned multiple cognitive distortion categories simultaneously. This formulation more accurately reflects clinical reality, where distortions co-occur and interact. Studies adopting multi-label setups typically model each cognitive distortion type as an independent binary label, which simplifies analysis and inter-annotator agreement calculations (Lybarger et al., 2022; Tauscher et al., 2023).
2.3 Auxiliary Tasks
Beyond isolated classification, some studies incorporate auxiliary tasks that extend the utility or interpretability of cognitive distortion models. These are not necessarily distinct problem categories, but rather downstream or complementary tasks that build upon classification outputs:
Reasoning generation.
This involves producing explanatory rationales for why a particular text span was classified as distorted. Methods such as Diagnosis of Thought (DoT) prompting (Chen et al., 2023) and the later ERD (Extraction, Reasoning, Debate) framework (Lim et al., 2024) aim to mimic clinician-like reasoning, improving model transparency and trustworthiness.
Reframing generation.
This focuses on producing healthier rephrasings of distorted thoughts, consistent with cognitive behavioural therapy interventions. Studies in this area (Sharma et al., 2023; Maddela et al., 2023) treat reframing as a natural extension of detection and classification.
Multi-task learning.
These setups combine cognitive distortion classification with related objectives, such as depression severity prediction (Wang et al., 2023a) or emotion cause extraction (Singh et al., 2023). These formulations tend to leverage the diagnostic value of cognitive distortion features to improve performance on auxiliary tasks.
Multi-modal approaches.
While the vast majority of computational cognitive distortion research focuses on textual data, recent work has begun to explore multi-modal approaches. For example, Singh et al. (2023) integrated text, audio, and video inputs from therapist-patient interactions to enhance cognitive distortion detection. Though still nascent, these efforts highlight the potential of multi-modal signals in capturing the subtle nuances of distorted thinking in real-world settings.
3 Taxonomies of cognitive distortions in natural language processing
Despite the shared objective of classifying cognitive distortions, studies seeking to label cognitive distortions have adopted diverse taxonomies. Without standardisation, it is hard to keep annotations consistent, compare models properly, or clearly interpret the results. Table 1 illustrates this fragmentation by mapping which cognitive distortion categories are recognised across the papers surveyed. While some categories, such as Labelling and Should Statements, are commonly used, others are inconsistently applied or redefined in various ways. Early conceptions of cognitive distortions were proposed by Beck (1963), who described patterns of dysfunctional thinking with examples such as arbitrary inference and overgeneralisation. This work was later popularised by Burns (1999), whose ten-category taxonomy is frequently cited in psychology literature [^2] While Burns (1999) is not the only taxonomy of cognitive distortions used in psychological contexts, and clinical practice often employs broader or alternative frameworks, our focus is on the inconsistency in how these definitions are interpreted and applied within computational research.. However, computational studies do not uniformly follow this framework.
Some papers, such as Shickel et al. (2020), draw on definitions from popular psychology sources (e.g., PsychCentral, Psychology Today), resulting in the inclusion of broader or differently framed categories. Other works make subtle assumptions in how the Burns taxonomy is applied, for example, by splitting Jumping to Conclusions into its subcategories Mind Reading and Fortune Telling, sometimes without explicit rationale.
Terminology also varies across studies. The cognitive distortion category referred to as All or Nothing Thinking also frequently appears under alternative labels such as Black and White Thinking, Polarised Thinking, or Dichotomous Reasoning. This terminological variety complicates efforts to harmonise and compare datasets. Since Rojas-Barahona et al. (2018) noted that CDs were ‘fairly well standardised’ in computational research, the field has grown considerably - from a handful of studies to 38 now in this survey - leading to a surge in differing taxonomies. The aim of Table 1 is to offer a consolidated view of this evolving landscape, providing a resource for future natural language processing research. Additionally, we include an appendix table (Table 3) listing synonyms, definitions, and hierarchical relationships between cognitive distortion types, to support more consistent and transparent classification efforts in computational contexts.
4 Datasets
Dataset Language Size (# Samples)* Labelling Annotators Access Literature Examples Wiemer-Hastings et al. (2004) English 261 Single-label (10) Expert Private Social Media Alhaj et al. (2022) Arabic 9,250 Single-label (5) Non-Expert (Unspecified) Private SocialCD-3K, Qi et al. (2024) Mandarin 3,407 Multi-label (12) Domain-Informed Public Aureus et al. (2021) English 586 Binary (2) Mixed Private Simms et al. (2017) English 459 Binary (2) Mixed Private Digital Mental Health Platform Rojas-Barahona et al. (2018) English 4,035 Multi-label (15) Expert Private Lin et al. (2024) Mandarin 4,001 Binary (2) Domain-Informed Public TherapistQA, Shreevastava and Foltz (2021) English 2,529 Multi-label (10) Non-Expert (Unspecified) Public MH-D, Shickel et al. (2020) English 1,799 Binary (2) Domain-Informed Private MH-C, Shickel et al. (2020) English 1,164 Single-label (15) Domain-Informed Private CBT-CD, Zhang et al. (2025) English 146 Multi-label (10) Expert Public Crowd-sourced Elsharawi and El Bolock (2024) English 34,370 Single-label (14) Expert Private PatternReframe, Maddela et al. (2023) English 9,688 Multi-label (10) Crowd-Generated Public CrowdDist, Shickel et al. (2020) English 7,666 Single-label (15) Crowd-Generated Private C2D2, Wang et al. (2023b) Mandarin 7,500 Single-label (7) Crowd-Generated Request Thinking Trap, Sharma et al. (2023) English 600 Multi-label (13) Expert Public Synthetic GPT-4 Synthetic, Babacan et al. (2025) English 2,000 Single-label (10) Automated (LLM) Public Clinical Intervention Lalk et al. (2024) German 104,557 Multi-label (14) Automated (Lexicon) Request Lybarger et al. (2022) English 7,436 Multi-label (5) Expert Private Hybrid (Mixed Domains) KoACD, Kim and Kim (2025) Korean 108,717 Single-label (10) Automated (LLM) Request GPT-4 Combined, Babacan et al. (2025) English 4,530 Single-label (10) Automated (LLM) Request CoDEC, Singh et al. (2023) English 3,773 Binary (2) Non-Expert (Unspecified) Request CoDeR, Singh et al. (2024) English 3,773 Binary (2) Trained Public Wang et al. (2023a) English 3,644 Single-label (11) Automated (BERT) Private Mostafa et al. (2021) English 2,409 Single-label (2) Domain-Informed Private
Table 2: Overview of datasets for cognitive distortion detection and classification, grouped by domain. See Appendix Table 4 for an expanded version with agreement metrics, access details, and subdomains. Corpus name, or earliest study to use it for cognitive distortion tasks. * Number of annotated units (e.g., posts, speech turns); for automated annotations, items processed. Number of cognitive distortion categories used, excluding ‘Undistorted’ for classification.
Datasets are the foundation for research on cognitive distortion classification, providing the labelled examples needed to develop and evaluate detection methods. However, existing datasets vary widely in scope, annotation practices, and accessibility. To organise this diversity, we group datasets by their underlying data sources and contexts of use.
4.1 Domains
We use the term domain to describe the broader context from which a dataset’s text data originates. Domains shape the linguistic style of examples, affect annotation reliability, and carry practical considerations such as data privacy and availability. The following six domains reflect the main sources of data in current cognitive distortion classification research.
Literature Examples.
Early work, such as Wiemer-Hastings et al. (2004), used cognitive distortion examples from existing psychological literature (Beck, 1979; Burns, 1999). While these examples are clear and well-labelled, they are typically idealised and explicit, limiting their applicability to real-world, patient-generated language.
Social Media Platforms.
Public posts from platforms such as Reddit (Aureus et al., 2021), Twitter (Alhaj et al., 2022), and Weibo (Qi et al., 2024) provide naturally occurring cognitive distortion instances in user-generated content. This domain offers large volumes of data but poses challenges related to linguistic noise and context ambiguity.
Digital Mental Health Platforms.
Peer-support services, such as Koko (Rojas-Barahona et al., 2018) and TaoConnect (Shickel et al., 2020), have been valuable sources of data rich in cognitive distortions. The widely used TherapistQA dataset (Shreevastava and Foltz, 2021) originates from a Kaggle Q&A repository and has since been extended in multiple studies (Chen et al., 2023; Babacan et al., 2025; Lalk et al., 2024).
Crowd-Sourced Approaches.
To tackle issues of data scarcity and privacy, several studies have turned to crowdworkers to generate or annotate cognitive distortion examples. Well-known corpora created this way include CrowdDist (Shickel et al., 2020), PatternReframe (Maddela et al., 2023), and Thinking Trap (Sharma et al., 2023). These datasets are scalable and flexible but often lack the subtlety of real-world data, as crowdworkers may produce overly explicit examples.
Clinical Interventions.
These datasets, derived from real therapeutic conversations, reflect how people communicate in real-world settings. Notable examples include annotated patient-therapist text message exchanges (Lybarger et al., 2022; Tauscher et al., 2023) and psychotherapy transcripts (Lalk et al., 2024). Multimodal corpora such as CoDEC and CoDeR (Singh et al., 2023, 2024) also fall into this category, although they involve a mix of authentic and staged interactions. Despite their value, these datasets are often subject to access restrictions due to privacy and ethical considerations.
Synthetic Datasets
Recent work has explored using large language models to generate synthetic cognitive distortion data. Babacan et al. (2025) created GPT-4-generated corpora, while Kim and Kim (2025) recently released KoACD, a Korean dataset augmenting social media data with synthetic samples. Synthetic datasets support balanced and scalable resource creation, but may fail to capture the nuanced linguistic and contextual patterns present in genuine human language.
4.2 Comparative Discussion
Each domain offers distinct strengths and drawbacks. Clinical intervention datasets are highly representative of real-world therapeutic contexts but are usually small and difficult to obtain. In contrast, social media and digital mental health platforms provide scalable, naturally occurring data, though they often exhibit linguistic noise and structural inconsistency. Crowd-sourced datasets allow for controlled creation of cognitive distortion examples but can introduce stylistic artefacts that may not mirror authentic language use. Synthetic datasets, including those generated by large language models, support large-scale experimentation and balanced class distributions, yet require thorough validation to ensure their realism. Most existing corpora are monolingual, predominantly in English, as summarised in Table 2. However, recent efforts have started expanding into other languages. For instance, Wang et al. (2023b) introduced C2D2, a Mandarin corpus, while Kim and Kim (2025) developed KoACD, a Korean dataset. Additional dataset details, including subdomains, access links, and inter-annotator agreement figures, are provided in Appendix Table 4.
4.3 Annotation Strategies
Annotation strategies vary considerably across datasets. Clinical and literature-derived corpora typically rely on expert annotators, prioritising label quality at the expense of scalability. In contrast, social media and crowd-sourced datasets often involve non-expert annotators, sometimes supported by brief, domain-specific training from qualified psychologists to improve consistency. While these strategies enable large-scale annotation, reliably classifying cognitive distortions remains challenging. The task demands subtle, often subjective judgements, and studies consistently report low inter-annotator agreement, particularly when annotators lack deeper domain expertise. To mitigate this, some works adopt strict inclusion criteria, only retaining examples where annotators fully agree on the label or a subset of distortion types (Aureus et al., 2021; Shickel et al., 2020). Though this approach improves label precision, it risks introducing bias by systematically excluding ambiguous or borderline cases - which are arguably the most reflective of real-world cognitive distortion occurrences. Balancing scalability, annotation quality, and representativeness therefore remains a core challenge in cognitive distortion dataset construction. Moreover, these annotation choices have important implications for how inter-annotator agreement is evaluated - a point we revisit (§6).
5 Modelling Approaches
This section outlines computational approaches to cognitive distortion detection and classification, grouped into six methodological categories that reflect major developments in the field. Our overview is informed by experimental findings from the surveyed literature, detailed in the supplementary materials.
5.1 Rule-Based
The first systems for cognitive distortion classification were rule-based, using hand-crafted keyword patterns and syntactic features. Wiemer-Hastings et al. (2004) developed COGNO, a system that mapped surface linguistic cues (e.g. verb tense, negation, person markers) to predefined cognitive distortion categories. It performed well on a 10-class single-label task (Macro-F1 = 0.61) but was only tested on ‘polished’ textbook-style cognitive distortion examples. One of the main strengths of rule-based systems is their interpretability, a feature still highly valued in clinical settings, where transparency is critical. For instance, Lalk et al. (2024) employed a manually curated list of n-grams, based on previous work (Bathina et al., 2021), to monitor distortion frequency in psychotherapy transcripts and predict patient depression severity.
5.2 Traditional Machine Learning (Feature-based)
As more annotated corpora became available, early rule-based systems gave way to feature-based statistical models. These approaches combined classic classifiers, such as logistic regression (LR) and support vector machines (SVMs), with engineered features such as Linguistic Inquiry and Word Count (LIWC) scores and Term Frequency-Inverse Document Frequency (TF-IDF) vectors. Simms et al. (2017) demonstrated that LR trained on LIWC features performed well on the detection task using Tumblr data, while Shickel et al. (2020) found that TF–IDF with LR outperformed CNNs on a 15-class single-label synthetic dataset (F1 = 0.68), indicating that shallow linguistic cues can remain competitive even in classification tasks. Similarly, Shreevastava and Foltz (2021) used SVM with smooth inverse frequency (SIF) embeddings, achieving strong performance (F1 = 0.77) on the detection task despite the insensitivity of SIF to word order.
5.3 Deep Learning with Static Embeddings
Static word embeddings such as Word2Vec (Mikolov et al., 2013) and GloVe (Pennington et al., 2014) introduced vector-based representations of words derived from co-occurrence patterns. These effectively captured word similarity but failed to account for contextual nuance (e.g. ‘riverbank’ vs ‘bank’). Despite this limitation, such embeddings formed the backbone of early deep-learning based cognitive distortion classifiers. For example, Rojas-Barahona et al. (2018) combined GloVe with CNNs for multi-label cognitive distortion classification, outperforming traditional models. Similarly, Mostafa et al. (2021) trained LSTM models using GloVe vectors to classify different cognitive distortion types, although their findings highlighted concerns around overfitting due to the use of limited and synthetic data. To address data sparsity in Arabic texts, Alhaj et al. (2022) applied contextual topic modelling (CTM), integrating static embeddings with domain-specific topic information via BERTopic - an approach that proved helpful in low-resource settings. While static embeddings offered richer representations than earlier methods, their inability to capture context, especially for polysemous words and subtle pragmatic distinctions, ultimately led to the rise of contextual models.
5.4 Transformer-based Architectures
Transformers, particularly BERT (Bidirectional Encoder Representations from Transformers) (Devlin et al., 2019), introduced contextual embeddings that capture word meaning based on the surrounding sentence, marking a significant leap forward in many natural language processing tasks, including cognitive distortion detection and classification. Shreevastava and Foltz (2021) used fine-tuned Sentence-BERT (SBERT) for binary cognitive distortion detection, showing notable improvements over earlier models. Similarly, Lybarger et al. (2022) demonstrated that incorporating conversational history further enhances classification performance on therapy dialogues. Domain-adapted transformers like MentalBERT (Ji et al., 2022), which are pre-trained on mental-health-related data, showed benefits over general-purpose models, while augmentation techniques such as mixup (Zhang et al., 2018) were explored to improve performance on rare cognitive distortion classes (minor improvements). Maddela et al. (2023) found that models like RoBERTa performed better than larger language models such as GPT-3.5 for this task, though distinguishing between closely related distortions remained a challenge. Overall, transformers brought significant gains in robustness and accuracy, but these came at the cost of reduced interpretability and a higher risk of overfitting, especially in datasets with extreme class imbalance. These limitations set the stage for the emergence of prompt-based models.
5.5 LLMs and Prompting Frameworks
large language models like GPT-3 have enabled cognitive distortion detection through natural language prompting, allowing models to perform the task without the need for tailored training. Recent work in this area can broadly be divided into two approaches: zero-shot prompting and chain-of-thought frameworks.
Zero-shot prompting.
Chen et al. (2023) introduced Diagnosis-of-Thought (DoT) prompting, guiding large language models to reason through cognitive distortion detection and classification with structured outputs. While this method sometimes outperformed fine-tuned transformers on the TherapistQA corpus, it also is prone to hallucinations and inconsistent rationales, especially on ambiguous cases. Similarly, Pico et al. (2025) compared multiple large language models, finding that well-prompted open-source models could approach the performance of larger proprietary models, though results were not always consistent across runs.
Chain-of-thought (CoT).
Lim et al. (2024) presented the ERD framework, where multiple large language model agents simulate therapist-like reasoning, extracting emotional cues and providing structured explanations. Though this method produced richer rationales, it was highly sensitive to how prompts were designed and faced challenges in scalability and validation. Another key application of large language models has been synthetic data generation. For example, Babacan et al. (2025) used GPT-4 to create a balanced cognitive distortion dataset. While initial results were promising, the synthetic data struggled to generalise to the noisier, more complex language found in real-world user inputs.
Overall, prompted large language models offer advantages in reducing training costs and improving interpretability, but they remain limited by issues such as prompt fragility, hallucinations, and inconsistent evaluation results.
5.6 Multimodal and Multi-task Architectures
To overcome the limitations of text-only models, recent research has explored incorporating additional modalities and joint tasks. Singh et al. (2023) introduced CoDEC, a multimodal dataset combining video, audio, and text from therapy simulations. By leveraging intonation and facial expressions, their model achieved improvements in detecting emotion-related cognitive distortions such as Emotional Reasoning. Building on CoDEC, Singh et al. (2024) developed CoDeR, which added annotated reasoning spans to support explanation-aware cognitive distortion classification. Multi-task learning has also been employed to utilise the diagnostic value of cognitive distortions. For example, Lee et al. (2021) repurposed micromodel outputs to improve depression and PTSD prediction, while Wang et al. (2023b) demonstrated that incorporating cognitive distortion frequency improved mental illness detection pipelines. These architectures have shown promise in boosting robustness, particularly in low-resource or noisy data settings. However, the scarcity of multimodal datasets and challenges around annotation and privacy continue to hinder wider adoption.
6 Evaluation
Despite progress in cognitive distortion classification, evaluation practices remain inconsistent, with studies differing markedly in their choice and reporting of metrics, which hinders comparability. Although F1 score is the most commonly used metric, distinctions between macro, micro, and weighted variants are frequently overlooked - a significant issue for class-imbalanced datasets where per-class performance is critical. Some studies now report AUPRC (Area Under the Precision–Recall Curve) to account for skewed label distributions, providing a more informative measure of performance on rare cognitive distortion types (Ding et al., 2022). Nonetheless, per-class metrics are still underreported, hiding weaknesses in addressing infrequent distortions. Similarly, dataset quality is often inconsistently assessed. Inter-annotator agreement (IAA) is either inconsistently measured or reported using incomparable metrics. While Cohen’s Kappa, which adjusts for chance, is typically more appropriate, many studies instead rely on raw agreement or pairwise F1, blurring the line between annotation reliability and model performance. Baseline comparisons are further complicated by inconsistent cognitive distortion taxonomies and datasets. We therefore recommend that researchers report per-class F1 scores, provide transparent IAA using chance-adjusted measures, and apply greater scrutiny to dataset origins.
7 Challenges & Future Directions
Despite recent progress, the automatic detection and classification of cognitive distortions remains a challenging task, both conceptually and computationally. In this section, we outline three key challenges currently limiting the field: (1) inconsistency in cognitive distortion taxonomies, (2) data scarcity and imbalance, and (3) the overreliance on short-form text. Addressing these issues is essential for improving model performance, evaluation fairness, and eventual clinical applicability.
7.1 Inconsistent cognitive distortion Taxonomies
A longstanding challenge is the lack of a standardised taxonomy for cognitive distortions. While foundational frameworks such as the Burns ten-category list (Burns, 1999) are commonly cited, computational studies diverge significantly in how they define, split, or rename distortion types. For instance, Jumping to Conclusions is frequently subdivided into Mind Reading and Fortune Telling, and terms such as All or Nothing Thinking appear under multiple aliases (e.g., Black and White Thinking, Polarised Thinking), making it difficult to compare models across studies, reproduce results, or interpret outputs reliably. These inconsistencies also affect annotation quality, as ambiguous or overly granular label sets introduce subjectivity and reduce inter-annotator agreement - an issue compounded by the lack of formal guidance on taxonomy use. While we provide a consolidated reference (Table 1), we do not claim it to be definitive. In the meantime, we encourage researchers to (i) report the source and rationale for their chosen taxonomy, (ii) avoid expanding taxonomies without empirical justification, and (iii) prioritise those grounded in clinical consensus where possible - with Burns’ taxonomy as a sensible default unless there is a well-founded reason to depart from it.
7.2 Data Scarcity, Imbalance, and Annotation Limits
The field remains constrained by a lack of large, high-quality datasets that capture authentic, context-rich examples of distorted thinking. Clinical corpora are scarce and often inaccessible due to privacy constraints, while many widely used datasets are synthetic, crowd-sourced, or compiled from multiple domains, and may lack the nuance and ambiguity of real-world language. This limits the complexity of distortions that models can learn and typically results in heavy class imbalance, with rare distortion types being underrepresented or excluded altogether. Although augmentation strategies such as mixup or back-translation offer minor gains for rare categories (Ding et al., 2022), a deeper issue lies in the ceiling imposed by annotation reliability itself. In some settings, even expert annotators show limited agreement, particularly for subtle or overlapping categories. For instance, Tauscher et al. (2023) report that for the presence of ‘Any Distortion’ in a text, human–human F1 agreement was 0.63, while a fine-tuned BERT model reached 0.62 - a very small difference. This suggests that for certain formulations, such as binary detection or high-frequency classes, current models may already be approaching the upper bound set by annotation quality. It also reinforces the need for clearer task definitions and more consistent annotation protocols before investing in model complexity. We recommend that future work (i) continues to report human–model comparison metrics where possible, (ii) prioritises the release of multilingual and multi-domain corpora, and (iii) explicitly documents how annotation disagreements are resolved. Without such efforts, further gains in model performance may reflect noise-fitting rather than genuine progress.
7.3 Overreliance on Short Text
The vast majority of existing datasets frame cognitive distortion detection at the sentence or single-utterance level. This simplifies annotation and model design but introduces strong limitations, as many distortions are context-dependent or only weakly signalled lexically. By stripping away discourse-level cues, models are forced to rely on surface-level patterns and may perform poorly on more ambiguous cases. Predictably, this has contributed to the dominance of distortion types with overt markers (e.g., Should Statements) in both datasets and model outputs. Empirical studies confirm the value of richer context: Lybarger et al. (2022) showed that including prior conversational turns improved detection F1 from 0.68 to 0.73, while frameworks like ERD (Lim et al., 2024) achieve greater interpretability by explicitly reasoning over multi-sentence inputs. Still, most benchmarks continue to prioritise short-form inputs. Moving forward, we argue that context-aware models should become the norm rather than the exception. In parallel, new datasets should prioritise multi-turn conversations, real patient narratives, and longer-form content that more closely mirrors therapeutic language.
8 Limitations
While this survey offers a structured overview of methods, datasets, and evaluation practices for cognitive distortion detection and classification in natural language processing, it has several limitations. The focus is primarily computational, with limited integration of insights from clinical psychology or cognitive science, and deeper conceptual analyses of cognitive distortions are beyond its scope. The survey also centres on English-language datasets and approaches, which may limit generalisability to other languages and cultural contexts. Although emerging work on multimodal and conversational systems is noted, the emphasis remains on text-based methods and classification tasks, rather than auxiliary tasks such as cognitive reframing. Finally, due to space constraints, some datasets and methods could not be covered in detail, and despite systematic efforts, some relevant studies may have been missed.
As with any literature survey, our analysis is constrained by the scope, reporting quality, and coverage of the included studies, and should be viewed as a snapshot of a rapidly evolving field. As such, some very recent preprints may not be included. While we have aimed for balanced representation, our synthesis reflects our methodological choices and interpretive framing, which may influence the emphasis placed on particular themes. To support transparency and future research, we plan to release our paper list.
9 Ethical Considerations
Given the nature of cognitive distortions within psychotherapy contexts, this survey acknowledges several important ethical considerations. As our work is a meta-analysis and synthesis of existing studies, we did not collect new data or propose new models. Nevertheless, the scope of our review touches on certain areas of concern that warrant attention.
Dataset Origins.
Many of the datasets discussed in this survey are derived from sources where individuals may have disclosed personal, and often highly sensitive, information. This is particularly true in domains such as digital mental health platforms, social media, and therapy transcripts. In reviewing these studies, we noted that some datasets have limited publicly available information on certain aspects of their origins or collection processes, often due to constraints inherited from upstream sources.
For instance, with TherapistQA, the authors provide clear documentation - including detailed labelling guidelines and procedures for resolving disagreements between annotators (Shreevastava and Foltz, 2021). However, because the dataset draws on an upstream public source, some provenance details reflect the level of information made available by that original source rather than any omission by the curators themselves. In this case, the publicly available version is based on a Kaggle Q&A dataset [^3] https://www.kaggle.com/datasets/arnmaud/therapist-qa/data, for which we have not found publicly accessible details specifying the original platform or data collection process. This situation is not unique to TherapistQA; several widely used mental health corpora draw on similar repositories, highlighting a broader and ongoing challenge in achieving complete transparency of data origins within the field.
Linguistic & Cultural Biases.
The literature we surveyed remains heavily focused on English-language data, with only limited, though encouragingly increasing, attention paid to other languages or cultural contexts. This linguistic bias introduces significant limitations, especially given that cognitive distortions are almost certainly shaped by cultural norms, stigma, and may manifest quite differently across populations. We repeat calls from prior work for the development and evaluation of more cognitive distortion classification methods that are sensitive to cross-linguistic and cross-cultural variation.
Risks of Misuse & Overreliance.
We also acknowledge that the automatic detection of cognitive distortions carries serious risks if applied irresponsibly, particularly outside of therapeutic settings. Misclassification or over-reliance on automated outputs could result in harm - reinforcing stigma, invalidating personal experiences, or leading to inappropriate interventions. We therefore stress that cognitive distortion classification systems should not be deployed without careful validation, the involvement of mental health professionals, and appropriate safeguards to protect user autonomy and well-being. As highlighted in Section 5, cognitive distortion classification performance is variable and often limited on rarer classes. As such, applications in clinical settings should be approached cautiously.
In presenting this survey, our aim is to support the natural language processing community in enabling more ethically considerate, transparent, and responsible research practices - particularly when working in sensitive domains such as mental health.
Acknowledgments
Some cognitive distortion examples in Table 3 were generated with the assistance of ChatGPT. The tool was also used in its intended role as an academic support resource to refine and polish author-written text. All AI-assisted content was carefully reviewed, verified, and edited by the authors.
This work was supported by the Engineering and Physical Sciences Research Council [grant number EP/W524475/1].
Appendix A Survey Methodology
To compile a comprehensive list of relevant research publications, we drew from the following sources:
  1. 1.
    Searches conducted across the ACL Anthology [^4] https://aclanthology.org/, arXiv [^5] https://arxiv.org/, PubMed [^6] https://pubmed.ncbi.nlm.nih.gov/, and IEEE Xplore [^7] https://ieeexplore.ieee.org/, with no date restrictions. Search queries included terms such as ‘cognitive distortion’ and ‘dysfunctional thought’.
  2. 2.
    Additional papers identified organically via Google Scholar, Semantic Scholar, and reference lists from relevant work.
After manual filtering, we retained 38 primary publications and preprints, spanning from 2004 to May 2025. Studies were included if they (i) implemented a cognitive distortion detection or classification model, (ii) introduced a cognitive distortion-related dataset, or (iii) computationally explored the taxonomy of cognitive distortions. To ensure reproducibility, we will publicly release all our search queries, filtering criteria, and the resulting paper list online upon publication.
Appendix B Psychological Foundations of cognitive distortions
This appendix offers an overview of cognitive distortions, providing additional context for natural language processing researchers who may be unfamiliar with them.
B.1 What are cognitive distortions?
As introduced in Section 1, cognitive distortions, sometimes called thinking errors, are habitual patterns of negatively biased or flawed thinking that shape how people interpret events, evaluate themselves, and respond to the world (Beck, 1963). There are many different types of cognitive distortions, as outlined in Table 3. In this paper, we use the term cognitive distortion taxonomy to refer to the particular set of cognitive distortions adopted in any given natural language processing study.
cognitive distortions are not restricted to clinical settings, as most people exhibit distorted thinking, often automatically, in response to certain situations (a key aspect studied in cognitive behavioural therapy). However, in conditions such as depression, anxiety, and post-traumatic stress disorder, these patterns tend to occur more frequently (Lalk et al., 2024), become harder to shift, and carry a heightened emotional impact.
B.2 Origins
The origins of cognitive distortions are often traced back to the early work of Beck in the 1960s. While he identified fewer types of distortions than are commonly recognised today, he did describe well-known types such as Overgeneralisation, as well as others that are less frequently cited. Quoted directly from his work (Beck, 1963) [^8] Both Arbitrary Interpretation and Selective Abstraction align reasonably well with the consolidated taxonomy reference in Table 1, corresponding to Jumping to Conclusions and Mental Filter, respectively.:
  • Arbitrary Interpretation – ‘the process of forming an interpretation of a situation, event, or experience when there is no factual evidence to support the conclusion, or when the conclusion is contrary to the evidence.’
  • Selective Abstraction – ‘focusing on a detail taken out of context, ignoring other more salient features of the situation, and conceptualising the whole experience on the basis of this element.’
The concept of cognitive distortions gained further traction in the 1980s through the work of David Burns, who outlined a widely used taxonomy of ten cognitive distortions in Feeling Good: The New Mood Therapy, later revising it in 1999 (Burns, 1999). Burns’ list more closely resembles the taxonomies commonly used in natural language processing studies today.
B.3 Context & Clinical Relevance
Although this survey focuses on computational approaches to cognitive distortion detection and classification, it is important to situate this area within its broader psychological context. Several therapeutic frameworks address distorted thinking, either directly or indirectly. For example:
Rational Emotive Behaviour Therapy (REBT)
where the focus is on the identification and challenge of irrational beliefs, which are broader than the discrete thought patterns typically considered as cognitive distortions. REBT focuses on core value beliefs (e.g., ‘I must be the best’) that tend to underlie many distortions, aiming to replace them with rational alternatives (Ellis, 1957, 1994; Dryden, 2021).
Schema therapy
moves past the present-moment focus of traditional cognitive behavioural therapy by addressing deep rooted, and often unhelpful, patterns, which are known as schemas. These can develop in childhood when core emotional needs are not met. Such schemas can lead to ongoing problems in how a person thinks, feels, behaves, and relates to others, often requiring longer term and more intensive treatment (Young et al., 2006).
Acceptance and Commitment Therapy (ACT)
does not frame problematic thinking in terms of cognitive distortions, but instead chooses to recognise unhelpful thoughts as a normal part of human thinking. The focus is not on challenging the content of these thoughts, but on changing the individual’s relationship to them through various strategies (Hayes et al., 2011).
While these approaches differ in focus, they generally agree on the importance of recognising distorted thinking patterns as a route to improved emotional or behavioural regulation. This clinical grounding remains a key motivation behind the computational modelling of cognitive distortions.
Appendix C Additional Tables
Code Cognitive Distortion Description Example Synonyms Burns’ Taxonomy Distortions (Burns, 1999) AON All or Nothing Thinking Viewing situations in black-and-white terms, without acknowledging nuance or grey areas. Since our method didn’t outperform all baselines in every metric, the entire study feels like a failure. Black and White Thinking, Polarised Thinking, Dichotomous Reasoning DQP Disqualifying the Positive Rejecting positive outcomes or feedback as unimportant, accidental, or unearned. Our paper was accepted, but probably only because the reviewers didn’t scrutinise it deeply enough. Discounting the Positive EMR Emotional Reasoning Believing that negative emotions reflect objective truths. I feel uneasy about presenting this model, so it must be inherently flawed in ways I’m not seeing. FTL Fortune Telling Predicting negative outcomes as inevitable, without sufficient evidence. Given how niche our contribution is, there’s no chance it will get noticed by the review committee. Negative Predictions, The Fortune Teller Error JTC Jumping to Conclusions Making assumptions with insufficient evidence. The editor’s brief reply likely means they’ve already decided to reject our manuscript. Jumping to Negative Conclusions LBL Labelling Defining oneself or others by a single trait or outcome. I misinterpreted that reviewer comment, clearly I’m not cut out for academic writing. Global Labelling, Labelling and Mislabelling MAG Magnification* Exaggerating the significance of errors or flaws. This small formatting mistake will probably make the reviewers think we lack attention to detail. Catastrophising* MIN Minimisation Downplaying the significance of positive outcomes, achievements, or strengths, reducing their perceived value or relevance. Sure, the paper was accepted, but it didn’t get the best reviews, so it doesn’t really count as a proper success. MTF Mental Filter Focusing exclusively on negative details. One weakness in our ablation study keeps bothering me, despite the overall positive experimental results. Filtering MDR Mind Reading Assuming you know what others are thinking, often negatively. The session chair looked disinterested, our work must have been irrelevant to the audience. OVG Overgeneralisation Drawing broad conclusions from a single incident. Since our last submission was desk-rejected, it’s obvious our current work will face the same fate. Overgeneralising PRS Personalisation Attributing external events or failures entirely to oneself. The collaboration didn’t materialise, probably because my proposal wasn’t convincing enough. Personalisation and Blame, Personalising, Blaming Oneself SHD Should Statements Holding rigid expectations about how oneself or others ought to behave. I should always produce novel ideas quickly, taking this long feels like professional incompetence. Shoulds, Inflexibility Other Distortions BRT Being Right Placing too high value on proving yourself correct, often at your own or others’ expense. I’m certain my annotation guidelines are the best. Any disagreement from the team simply indicates they don’t understand the task properly. Always Being Right BLM Blaming Attributing too high responsibility for negative outcomes to others, avoiding self-reflection or your own shared responsibility. The demo crashed because the organisers didn’t provide adequate technical support, not because of any oversight on our side. Blaming Others CAT Catastrophising* Imagining worst-case scenarios and exaggerating potential negative consequences far beyond their realistic likelihood. If this preprint has a minor oversight, it could irreparably damage our lab’s reputation and future collaborations. CMP Comparing Measuring self-worth against others in a way that undermines your own accomplishments. Another lab published a similar paper first - clearly they’re much more capable researchers than we are. Comparing and Despairing, Comparison CTL Control Fallacy Believing either complete control over everything or total helplessness in a situation, without middle ground. If I don’t oversee every single preprocessing step myself, the entire pipeline will end up flawed. FOC Fallacy of Change Assuming others should or will change to meet your own personal expectations. If only the dataset creators had annotated according to our taxonomy, our analysis would be so much clearer. Control of Fallacies FOF Fallacy of Fairness Presuming life or systems must work in a way that aligns with personal standards of fairness. It’s unfair that methodologically weaker papers receive more attention just because they’re trendy. HRF Heaven’s Reward Fallacy Expecting a guaranteed reward for one’s hard work. After months of hyperparameter tuning, this model surely deserves to be the new state-of-the-art. LFT Low Frustration Tolerance* Overestimating the severity of minor inconveniences. Dealing with this reviewer rebuttal feels impossible. I can’t imagine going through it again. NFE Negative Feeling or Emotion Taking emotional discomfort as proof something is wrong. Feeling stuck while writing this paper draft surely means the research itself is inherently flawed.
Table 3: Categories of cognitive distortions observed in computational research. Descriptions and examples are reflective of common interpretations of these distortions in natural language processing contexts. All examples are fictional and not about any specific work or group. To ensure consistency across studies, we also include synonyms and related terms where applicable. Jumping to Conclusions (JTC) is frequently considered a parent category that includes Fortune Telling (FTL) and Mind Reading (MDR). * Although Magnification (MAG) and Catastrophising (CAT) are often treated as equivalent, we list them separately to highlight subtle conceptual distinctions, following prior work (Lalk et al., 2024; Agarwal and Sirts, 2025). Similarly, Low Frustration Tolerance (LFT), while similar to CAT, is presented as a distinct category.
Dataset Language Subdomain Size (# Samples)* Labelling Annotators Agreement Access Literature Examples Wiemer-Hastings et al. (2004) English Psychology literature 261 Single-label (10) Expert Private Social Media Alhaj et al. (2022) Arabic Twitter 9,250 Single-label (5) Non-Expert (Unspecified) $\kappa=0.817_{c}$ Private SocialCD-3K, Qi et al. (2024) Mandarin Weibo ‘Zoufan’ blog 3,407 Multi-label (12) Domain-Informed Public1 Aureus et al. (2021) English Reddit: r/COVID19_support 586 Binary (2) Mixed Private Simms et al. (2017) English Tumblr 459 Binary (2) Mixed Private Digital Mental Health Platform Rojas-Barahona et al. (2018) English Koko 4,035 Multi-label (15) Expert $\kappa=0.61_{c}$ Private2 Lin et al. (2024) Mandarin PsyQA counselling forums 4,001 Binary (2) Domain-Informed $\mathrm{JP}=0.88_{d}$ Public3 TherapistQA, Shreevastava and Foltz (2021) English 2,529 Multi-label (10) Non-Expert (Unspecified) $\mathrm{JP}=0.34_{c},0.61_{d}$ Public4 MH-D, Shickel et al. (2020) English TaoConnect 1,799 Binary (2) Domain-Informed Private MH-C, Shickel et al. (2020) English TaoConnect 1,164 Single-label (15) Domain-Informed Private CBT-CD, Zhang et al. (2025) English Patient-therapist QA 146 Multi-label (10) Expert Public5 Crowd-sourced Elsharawi and El Bolock (2024) English 34,370 Single-label (14) Expert Private PatternReframe, Maddela et al. (2023) English MTurk, Mephisto 9,688 Multi-label (10) Crowd-Generated $\alpha=0.355_{c}$ Public6 CrowdDist, Shickel et al. (2020) English MTurk 7,666 Single-label (15) Crowd-Generated Private C2D2, Wang et al. (2023b) Mandarin 7,500 Single-label (7) Crowd-Generated $\kappa=0.67_{c}$ Request7 Thinking Trap, Sharma et al. (2023) English 600 Multi-label (13) Expert Public8 Synthetic GPT-4 Synthetic, Babacan et al. (2025) English GPT-4 2,000 Single-label (10) Automated (LLM) Public9 Clinical Intervention Lalk et al. (2024) German CBT psychotherapy transcripts 104,557 Multi-label (14) Automated (Lexicon) Request10 Lybarger et al. (2022) English Patient-therapist text exchanges 7,436 Multi-label (5) Expert $\kappa=0.53_{d}$ Private Hybrid (Mixed Domains) KoACD, Kim and Kim (2025) Korean NAVER Knowledge iN + LLM 108,717 Single-label (10) Automated (LLM) $\kappa=0.78$ Request11 GPT-4 Combined, Babacan et al. (2025) English GPT-4 synthetic + TherapistQA 4,530 Single-label (10) Automated (LLM) Request12 CoDEC, Singh et al. (2023) English Real + staged patient-therapist videos 3,773 Binary (2) Non-Expert (Unspecified) $F=0.83_{d}$ Request13 CoDeR, Singh et al. (2024) English Real + staged patient-therapist videos 3,773 Binary (2) Trained $F=0.83_{d}$ Public14 Wang et al. (2023a) English Lit. examples + social media augment 3,644 Single-label (11) Automated (BERT) Private Mostafa et al. (2021) English Twitter, Surveys, HappyDB 2,409 Single-label (2) Domain-Informed Private 1 https://github.com/HongzhiQ/SupervisedVsLLM-EfficacyEval/tree/main/data/SocialCD-3k 2 https://github.com/YinpeiDai/NAUM 3 https://github.com/405200144/Dataset-of-Cognitive-Distortion-detection-and-Positive-Reconstruction/tree/main 4 https://www.kaggle.com/datasets/sagarikashreevastava/cognitive-distortion-detetction-dataset 5 https://huggingface.co/datasets/Psychotherapy-LLM/CBT-Bench 6 https://github.com/facebookresearch/ParlAI/tree/main/projects/reframe_thoughts 7 https://github.com/bcwangavailable/C2D2-Cognitive-Distortion 8 https://github.com/behavioral-data/Cognitive-Reframing 9 https://huggingface.co/datasets/halilbabacan/cognitive_distortions_gpt4 10 https://osf.io/rsy4z/?view_only=41dc962f0c924c0e87e7bfc044535bd3 11 https://github.com/cocoboldongle/KoACD 12 https://huggingface.co/datasets/halilbabacan/combined_synthetic_cognitive_distortions 13 https://www.iitp.ac.in/˜ai-nlp-ml/resources.html#DeCoDE-CoDEC 14 https://github.com/clang1234/ZS-CoDR.git
Table 4: Extended overview of datasets for cognitive distortion detection and classification, grouped by domain. Agreement metrics: $\kappa$ = Cohen’s kappa; $\alpha$ = Krippendorff’s alpha; $F$ = Fleiss’s kappa; JP = joint probability; d = detection; c = classification. ‘–’ indicates not applicable or not reported. Corpus name, or earliest study to use it for cognitive distortion tasks. * Number of annotated units (e.g., posts, speech turns); for automated methods, items processed. Number of cognitive distortion categories used, excluding ‘Undistorted’ for classification.

### v2

@@
 Abstract
-As interest grows in the application of natural language processing techniques to mental health, a growing body of work explores the automatic detection and classification of cognitive distortions.
-cognitive distortions are habitual patterns of negatively biased or flawed thinking that distort how people perceive events, judge themselves, and react to the world around them.
-Identifying and addressing them is an important part of therapy.
-Despite its momentum, the field remains fragmented, with inconsistencies in cognitive distortion taxonomies, task formulations, and evaluation practices.
-This survey reviews 38 studies spanning two decades, providing a structured overview of datasets, modelling approaches, and evaluation strategies.
-We provide a consolidated cognitive distortion taxonomy reference, summarise common task setups, and highlight open challenges to support more coherent and reproducible research in this emerging area.
+As interest grows in applying natural language processing techniques to mental health, an expanding body of work explores the automatic detection and classification of cognitive distortions.
+cognitive distortions are habitual patterns of negatively biased or flawed thinking that distort how people perceive events, judge themselves, and react to the world.
+Identifying and addressing them is a central goal of therapy.
+Despite this momentum, the field remains fragmented, with inconsistencies in cognitive distortion taxonomies, task formulations, and evaluation practices limiting comparability across studies.
+This survey presents the first comprehensive review of 38 studies spanning two decades, mapping how cognitive distortions have been implemented in computational research and evaluating the methods applied.
+We provide a consolidated cognitive distortion taxonomy reference, summarise common task setups, and highlight persistent challenges to support more coherent and reproducible research.
+Alongside our review, we introduce practical resources, including curated evaluation metrics from surveyed papers, a standardised datasheet template, and an ethics flowchart, available online.
+[^1] https://github.com/archiesage/cognitive-distortion-nlp-survey
 A Survey of Cognitive Distortion Detection and Classification in NLP
@@
 These distortions play a central role in emotional distress and are a core target of evidence-based psychological interventions such as cognitive behavioural therapy (Beck, 1963; Burns, 1999).
-Common examples include Catastrophising (‘I let them down, so they’ll never trust me again’), Mind Reading (‘They haven’t replied, so they must be angry at me’), and All or Nothing Thinking (‘If I don’t get this right the first time, I’m a complete failure’).
-[^1] A complete list of cognitive distortion definitions, examples, and synonyms is provided in Table 3, with additional psychological background in Appendix B.
-These patterns often appear intuitive or harmless at first glance, but they have been shown to maintain and exacerbate conditions like depression, anxiety, and post-traumatic stress disorder.
+Common examples [^2] A complete list of cognitive distortion definitions, examples, and synonyms is provided in Table 3, with additional psychological background in Appendix B.
+include Catastrophising (‘I let them down, so they’ll never trust me again’), Mind Reading (‘They haven’t replied, so they must be angry at me’), and All or Nothing Thinking (‘If I don’t get this right the first time, I’m a complete failure’).
+These patterns often appear intuitive or harmless at first, but they have been shown to maintain and exacerbate conditions like depression, anxiety, and post-traumatic stress disorder.
 In therapeutic settings, recognising and reframing such distortions is a core goal of cognitive behavioural therapy.
@@
 (ii) It defines and analyses task setups, datasets, computational methods, and approaches to performance evaluation, highlighting key patterns and gaps (§2, §4, §5, §6).
-(iii) It identifies open challenges and proposes concrete directions to guide future research (§7).
+(iii) It identifies open challenges (§7) and proposes best practices to guide future research (§8).
 In doing so, this paper aims to enable more consistent, comparable, and reproducible work in this emerging field.
@@
 Early conceptions of cognitive distortions were proposed by Beck (1963), who described patterns of dysfunctional thinking with examples such as arbitrary inference and overgeneralisation.
-This work was later popularised by Burns (1999), whose ten-category taxonomy is frequently cited in psychology literature [^2] While Burns (1999) is not the only taxonomy of cognitive distortions used in psychological contexts, and clinical practice often employs broader or alternative frameworks, our focus is on the inconsistency in how these definitions are interpreted and applied within computational research..
+This work was later popularised by Burns (1999), whose ten-category taxonomy is frequently cited in psychology literature.
+[^3] While Burns (1999) is not the only taxonomy of cognitive distortions used in psychological contexts, and clinical practice often employs broader or alternative frameworks, our focus is on the inconsistency in how these definitions are interpreted and applied within computational research.
 However, computational studies do not uniformly follow this framework.
@@
 Though this approach improves label precision, it risks introducing bias by systematically excluding ambiguous or borderline cases - which are arguably the most reflective of real-world cognitive distortion occurrences.
-Balancing scalability, annotation quality, and representativeness therefore remains a core challenge in cognitive distortion dataset construction.
-Moreover, these annotation choices have important implications for how inter-annotator agreement is evaluated - a point we revisit (§6).
 5 Modelling Approaches
 This section outlines computational approaches to cognitive distortion detection and classification, grouped into six methodological categories that reflect major developments in the field.
-Our overview is informed by experimental findings from the surveyed literature, detailed in the supplementary materials.
 5.1 Rule-Based
@@
 (2017) demonstrated that LR trained on LIWC features performed well on the detection task using Tumblr data, while Shickel et al.
-(2020) found that TF–IDF with LR outperformed CNNs on a 15-class single-label synthetic dataset (F1 = 0.68), indicating that shallow linguistic cues can remain competitive even in classification tasks.
+(2020) found that TF-IDF with LR outperformed CNNs on a 15-class single-label synthetic dataset (F1 = 0.68), indicating that shallow linguistic cues can remain competitive even in classification tasks.
 Similarly, Shreevastava and Foltz (2021) used SVM with smooth inverse frequency (SIF) embeddings, achieving strong performance (F1 = 0.77) on the detection task despite the insensitivity of SIF to word order.
@@
 Static word embeddings such as Word2Vec (Mikolov et al., 2013) and GloVe (Pennington et al., 2014) introduced vector-based representations of words derived from co-occurrence patterns.
-These effectively captured word similarity but failed to account for contextual nuance (e.g.
-‘riverbank’ vs ‘bank’).
+These effectively captured word similarity but failed to account for contextual nuance (e.g., ‘riverbank’ vs ‘bank’).
 Despite this limitation, such embeddings formed the backbone of early deep-learning based cognitive distortion classifiers.
@@
 Similarly, Lybarger et al.
-(2022) demonstrated that incorporating conversational history further enhances classification performance on therapy dialogues.
+(2022) demonstrated that incorporating conversational history further improves classification performance on therapy dialogues.
 Domain-adapted transformers like MentalBERT (Ji et al., 2022), which are pre-trained on mental-health-related data, showed benefits over general-purpose models, while augmentation techniques such as mixup (Zhang et al., 2018) were explored to improve performance on rare cognitive distortion classes (minor improvements).
@@
 However, the scarcity of multimodal datasets and challenges around annotation and privacy continue to hinder wider adoption.
+5.7 Feasibility of Meta-Analysis
+While we initially intended to include comparative performance tables, meaningful aggregation proved infeasible.
+The studies diverge across critical axes - (i) task formulation, (ii) cognitive distortion taxonomy choice, (iii) dataset domain, (iv) evaluation metrics, (v) granularity of the unit of analysis, (vi) context inclusion and window size, and (vii) modality - often incompatibly.
+Sample sizes within aligned subgroups are too small for robust comparison, and inter-annotator agreement is reported inconsistently, further limiting comparability.
+Pooled tables would therefore risk suggesting misleading trends.
+We release the full set of extracted results in our GitHub repository, enabling researchers to build their own comparisons.
+In this paper, we restrict ourselves to qualitative synthesis, deferring formal meta-analysis until the evidence base is larger and more standardised.
 6 Evaluation
@@
 Although F1 score is the most commonly used metric, distinctions between macro, micro, and weighted variants are frequently overlooked - a significant issue for class-imbalanced datasets where per-class performance is critical.
-Some studies now report AUPRC (Area Under the Precision–Recall Curve) to account for skewed label distributions, providing a more informative measure of performance on rare cognitive distortion types (Ding et al., 2022).
+Some studies now report AUPRC (Area Under the Precision-Recall Curve) to account for skewed label distributions, providing a more informative measure of performance on rare cognitive distortion types (Ding et al., 2022).
 Nonetheless, per-class metrics are still underreported, hiding weaknesses in addressing infrequent distortions.
 Similarly, dataset quality is often inconsistently assessed.
-Inter-annotator agreement (IAA) is either inconsistently measured or reported using incomparable metrics.
-While Cohen’s Kappa, which adjusts for chance, is typically more appropriate, many studies instead rely on raw agreement or pairwise F1, blurring the line between annotation reliability and model performance.
+Inter-annotator agreement is either inconsistently measured or reported using incomparable metrics.
+While Cohen’s Kappa ($\kappa$), which adjusts for chance, is typically more appropriate, many studies instead rely on raw agreement or non-standard metrics, blurring the line between annotation reliability and model performance.
 Baseline comparisons are further complicated by inconsistent cognitive distortion taxonomies and datasets.
-We therefore recommend that researchers report per-class F1 scores, provide transparent IAA using chance-adjusted measures, and apply greater scrutiny to dataset origins.
 7 Challenges & Future Directions
@@
 These inconsistencies also affect annotation quality, as ambiguous or overly granular label sets introduce subjectivity and reduce inter-annotator agreement - an issue compounded by the lack of formal guidance on taxonomy use.
-While we provide a consolidated reference (Table 1), we do not claim it to be definitive.
-In the meantime, we encourage researchers to (i) report the source and rationale for their chosen taxonomy, (ii) avoid expanding taxonomies without empirical justification, and (iii) prioritise those grounded in clinical consensus where possible - with Burns’ taxonomy as a sensible default unless there is a well-founded reason to depart from it.
 7.2 Data Scarcity, Imbalance, and Annotation Limits
 The field remains constrained by a lack of large, high-quality datasets that capture authentic, context-rich examples of distorted thinking.
-Clinical corpora are scarce and often inaccessible due to privacy constraints, while many widely used datasets are synthetic, crowd-sourced, or compiled from multiple domains, and may lack the nuance and ambiguity of real-world language.
+Clinical corpora are scarce and often inaccessible due to privacy constraints, while many widely used datasets are synthetic, crowd-sourced, or compiled from multiple domains, which may lack the nuance and ambiguity of real-world language.
 This limits the complexity of distortions that models can learn and typically results in heavy class imbalance, with rare distortion types being underrepresented or excluded altogether.
@@
 For instance, Tauscher et al.
-(2023) report that for the presence of ‘Any Distortion’ in a text, human–human F1 agreement was 0.63, while a fine-tuned BERT model reached 0.62 - a very small difference.
+(2023) report that for the presence of ‘Any Distortion’ in a text, human F1 agreement was 0.63, while a fine-tuned BERT model reached 0.62 - a very small difference.
 This suggests that for certain formulations, such as binary detection or high-frequency classes, current models may already be approaching the upper bound set by annotation quality.
 It also reinforces the need for clearer task definitions and more consistent annotation protocols before investing in model complexity.
-We recommend that future work (i) continues to report human–model comparison metrics where possible, (ii) prioritises the release of multilingual and multi-domain corpora, and (iii) explicitly documents how annotation disagreements are resolved.
-Without such efforts, further gains in model performance may reflect noise-fitting rather than genuine progress.
 7.3 Overreliance on Short Text
@@
 Predictably, this has contributed to the dominance of distortion types with overt markers (e.g., Should Statements) in both datasets and model outputs.
-Empirical studies confirm the value of richer context: Lybarger et al.
-(2022) showed that including prior conversational turns improved detection F1 from 0.68 to 0.73, while frameworks like ERD (Lim et al., 2024) achieve greater interpretability by explicitly reasoning over multi-sentence inputs.
+Empirical studies confirm the value of richer context, showing that including prior conversational turns improved detection F1 from 0.68 to 0.73 (Lybarger et al., 2022), while frameworks such as ERD achieve greater interpretability by explicitly reasoning over multi-sentence inputs (Lim et al., 2024).
 Still, most benchmarks continue to prioritise short-form inputs.
@@
 In parallel, new datasets should prioritise multi-turn conversations, real patient narratives, and longer-form content that more closely mirrors therapeutic language.
+8 Best Practices and Recommendations
+Alongside consolidating existing research, it is important to address the main sources of fragmentation in the field.
+We therefore propose a set of best practices for future work, which should be followed where possible or clearly justified if deviated from.
+8.1 Taxonomy Adoption
+As shown in Table 1, inconsistent use of cognitive distortion taxonomies has made cross-study comparison difficult.
+In the absence of a universally accepted taxonomy, we recommend Burns’ taxonomy [^4] We provide recommended label sets online.
+(Burns, 1999) as a sensible default, since it is the most widely cited and most alternatives used in natural language processing are partial reinterpretations of it.
+Researchers should (i) report the source and rationale for their chosen taxonomy, (ii) avoid introducing new or expanded taxonomies without justification, and (iii) prioritise taxonomies grounded in clinical consensus.
+Reliance on loosely defined online taxonomies is discouraged.
+Where deviations from Burns’ taxonomy are necessary, the rationale should be documented in the study, or, in the case of new datasets, in the corresponding datasheet, for which we provide a template online.
+8.2 Unambiguous Evaluation Reporting
+Inconsistent reporting is a major barrier to comparing results across studies.
+To improve comparability, we recommend that future work (i) clearly state the task formulation (detection, single-label, or multi-label classification; §2), (ii) specify the analysis unit (sentence, turn, session) and, if relevant, document the exact context window, (iii) indicate the cognitive distortion taxonomy used, (iv) report per-class scores alongside macro and weighted F1555For imbalanced or multi-label settings, AUPRC (micro or macro) and per-class PR curves may be more informative.
+Ultimately, metrics should fit the use case.
+with unambiguous labels, and (v) explain the choice of evaluation metrics.
+While a single metric cannot suit all applications, departures from macro or weighted F1 should always be accompanied by a clear rationale.
+8.3 Dataset Development and Use
+For researchers creating new cognitive distortion datasets, we recommend providing a datasheet that documents the dataset’s origin, annotation protocol, size, taxonomy, analysis unit, and licensing.
+To support this, we provide a standardised datasheet template in our GitHub repository.
+For reuse of existing datasets, we encourage researchers to apply our Ethics Flowchart, which provides practical guidance on assessing provenance, consent, and documentation before experimentation.
+8.4 Annotation Reliability and Inter-annotator Agreement
+Annotation processes should be reported transparently.
+Researchers should provide standard inter-annotator agreement metrics (e.g., Cohen’s $\kappa$, Fleiss’ $\kappa$, Krippendorff’s $\alpha$) rather than vague statements of ‘agreement’.
+The rationale for the chosen metric should be stated, and partial dataset sampling for inter-annotator agreement is acceptable provided that procedures are clearly documented.
+We further recommend that future work (i) reports human-model performance comparison metrics where possible, (ii) prioritises the release of multilingual and multi-domain corpora, and (iii) explicitly documents how annotation disagreements are resolved.
+Without such practices, gains in model performance may reflect noise-fitting rather than genuine progress.
+8.5 Code and Dataset Release
+To support replication and benchmarking, researchers should release code and, where licensing and privacy considerations permit, datasets.
+As shown in Table 4, many of the surveyed studies do not provide public implementations.
+We recommend that future work adopt code and data release as standard practice, in line with recent broader calls for stronger reproducibility standards in AI research and governance (Semmelrock et al., 2025; Mason-Williams and Mason-Williams, 2025).
 9 Limitations
@@
 While we have aimed for balanced representation, our synthesis reflects our methodological choices and interpretive framing, which may influence the emphasis placed on particular themes.
-To support transparency and future research, we plan to release our paper list.
 10 Ethical Considerations
 Given the nature of cognitive distortions within psychotherapy contexts, this survey acknowledges several important ethical considerations.
-As our work is a meta-analysis and synthesis of existing studies, we did not collect new data or propose new models.
+As our work is a synthesis of existing studies, we did not collect new data or propose new models.
 Nevertheless, the scope of our review touches on certain areas of concern that warrant attention.
@@
 However, because the dataset draws on an upstream public source, some provenance details reflect the level of information made available by that original source rather than any omission by the curators themselves.
-In this case, the publicly available version is based on a Kaggle Q&A dataset [^3] https://www.kaggle.com/datasets/arnmaud/therapist-qa/data, for which we have not found publicly accessible details specifying the original platform or data collection process.
+In this case, the publicly available version is based on a Kaggle Q&A dataset, [^6] https://www.kaggle.com/datasets/arnmaud/therapist-qa/data for which we have not found publicly accessible details specifying the original platform or data collection process.
 This situation is not unique to TherapistQA; several widely used mental health corpora draw on similar repositories, highlighting a broader and ongoing challenge in achieving complete transparency of data origins within the field.
@@
 Acknowledgments
-Some cognitive distortion examples in Table 3 were generated with the assistance of ChatGPT.
-The tool was also used in its intended role as an academic support resource to refine and polish author-written text.
-All AI-assisted content was carefully reviewed, verified, and edited by the authors.
-This work was supported by the Engineering and Physical Sciences Research Council [grant number EP/W524475/1].
+This work was supported by the Engineering and Physical Sciences Research Council [grant number EP/W524475/1]. We thank the anonymous reviewers for their constructive feedback.
 Appendix A Survey Methodology
@@
 Studies were included if they (i) implemented a cognitive distortion detection or classification model, (ii) introduced a cognitive distortion-related dataset, or (iii) computationally explored the taxonomy of cognitive distortions.
-To ensure reproducibility, we will publicly release all our search queries, filtering criteria, and the resulting paper list online upon publication.
+To support reproducibility, we release our supplementary resources and paper list, along with corrections and updates, on GitHub: https://github.com/archiesage/cognitive-distortion-nlp-survey.
 Appendix B Psychological Foundations of cognitive distortions
@@
 While he identified fewer types of distortions than are commonly recognised today, he did describe well-known types such as Overgeneralisation, as well as others that are less frequently cited.
-Quoted directly from his work (Beck, 1963) [^8] Both Arbitrary Interpretation and Selective Abstraction align reasonably well with the consolidated taxonomy reference in Table 1, corresponding to Jumping to Conclusions and Mental Filter, respectively.:
-  • Arbitrary Interpretation – ‘the process of forming an interpretation of a situation, event, or experience when there is no factual evidence to support the conclusion, or when the conclusion is contrary to the evidence.’
-  • Selective Abstraction – ‘focusing on a detail taken out of context, ignoring other more salient features of the situation, and conceptualising the whole experience on the basis of this element.’
+Quoted directly from his work (Beck, 1963): [^11] Both Arbitrary Interpretation and Selective Abstraction align reasonably well with the consolidated taxonomy reference in Table 1, corresponding to Jumping to Conclusions and Mental Filter, respectively.
+  • Arbitrary Interpretation - ‘the process of forming an interpretation of a situation, event, or experience when there is no factual evidence to support the conclusion, or when the conclusion is contrary to the evidence.’
+  • Selective Abstraction - ‘focusing on a detail taken out of context, ignoring other more salient features of the situation, and conceptualising the whole experience on the basis of this element.’
 The concept of cognitive distortions gained further traction in the 1980s through the work of David Burns, who outlined a widely used taxonomy of ten cognitive distortions in Feeling Good: The New Mood Therapy, later revising it in 1999 (Burns, 1999).